OpenStudy (anonymous):

Is (-1,5) a solutiom of this system y= -x + 4 y=-1x (that is a fraction) 5 Please give me procedure Help :)

OpenStudy (whpalmer4):

(-1,5) is x = -1, y = 5. Do the equations work with those values? If they do, then it is a solution. If they don't, it isn't.

OpenStudy (anonymous):

So\[y=-x+4\]\[y=\frac{-x}{5}\]If you look at the second equation, you can see that it is already solved for y. So in the first equation, wherever there is y you can replace it with -x/5. So:\[\frac{-x}{5}=-x+4\]\[-x=-5x+20\]\[4x=20\]\[x=5\]Then you can use the second equation to find out what y is.

OpenStudy (anonymous):

thanks :)

OpenStudy (whpalmer4):

Graphically, you have two lines, y = -x + 4 (slopes 1:1 from upper left to lower right, crosses y-axis at 4), and y = -x/5 (slopes 5:1 from upper left to lower right, crosses y-axis at 0). The solution to the system is the point at which they cross (which they do as they are not parallel). The $64,000 is do they cross at (-1,5)?

OpenStudy (whpalmer4):

That's "The $64,000 question"...

OpenStudy (whpalmer4):

And that should be slopes 1:5...sigh...

OpenStudy (whpalmer4):

Looking at the graph, (-1,5) isn't a solution, but (5,-1) looks like it might be.

OpenStudy (whpalmer4):

Sketching a quick graph is often a good check to see if your answer makes sense if you tend to make mistakes in your algebra.